11) A man on top of a building 30 m high drops a ball. At the same time a man on the ground directly beneath the ball fires a gun. What initial velocity must the bullet have in order to hit the ball 2 m below the top of the building?
1
Expert's answer
2020-10-02T07:15:14-0400


In time "t" a free falling ball covers a distance "h":


"h = v_0t + \dfrac{gt^2}{2}"

where "g = 9.81m\/s^2" is the gravitational acceleration and "v_0 = 0" (no starting velocity).

Thus, the ball will cover "h_1 = 2m" in the following time:


"t_{meet} = \sqrt{\dfrac{2h_1}{g}}"

The distance covered by the bullet in time "t" is:


"h = v_0t - \dfrac{gt^2}{2}"

where "v_0" is bullet's initial velocity. In time "t_{meet}" the bullet should cover the distnace "h_2". Thus:


"h_2 = v_0t_{meet} - \dfrac{gt_{meet}^2}{2}"


Then, the initial velocity is:


"v_0 = \dfrac{h_2}{t_{meet}} + \dfrac{gt_{meet}}{2}"

Substituting "t_{meet}", obtain:

"v_0 = h_2\sqrt{\dfrac{g}{2h_1}} + \sqrt{\dfrac{gh_1}{2}}"

In numbers:


"v_0 = 28\cdot \sqrt{\dfrac{9.81}{2\cdot 2}} + \sqrt{\dfrac{9.81\cdot 2}{2}}\approx 46.98m/s"

Answer. 46.98 m/s.